autumn has come and you need to empty your swimming pool at home. you have an older water pump that can empty your swimming pool in 5 hours. you also just bought a new water pump since it was on sale and it can empty your swimming pool in 3 hours. you want to get the pool drained as quickly as possible so you can go out for a run. you decide to use both water pumps to empty your swimming pool at once. how long will it take the pool using both water pumps?
Answer by stanbon(75887)

autumn has come and you need to empty your swimming pool at home. you have an older water pump that can empty your swimming pool in 5 hours. you also just bought a new water pump since it was on sale and it can empty your swimming pool in 3 hours. you want to get the pool drained as quickly as possible so you can go out for a run. you decide to use both water pumps to empty your swimming pool at once. how long will it take the pool using both water pumps?
---
Older pump DATA:
time = 5 hrs/job ; rate = 1/5 job/hr
----------------
Newer pump DATA:
time = 3 hrs/job ; rate = 1/3 job/hr
-----------------
Together DATA:
time = x hrs/job ; rate = 1/x job/hr
====================================
Equation:
rate + rate = together rate
1/5 + 1/3 = 1/x
---
Multiply thru by 15x to get:
3x + 5x = 15
8x = 15
x = 15/8 hrs (time required for both pumpts to empty the pool together)
========================
Cheers,
Stan H.